Suppose that an element has two isotopes. The average atomic mass of the element is  49.237 u. One isotope has a mass of 41.834 u

 and an abundance of 35.94%.

Calculate the mass of the second isotope.


1
Expert's answer
2021-02-22T06:41:29-0500

Solution:

The average atomic mass of an element is the sum of the masses of its isotopes, each multiplied by its natural abundance.

Average atomic mass = f1M1 + f2M2 +… + fnMn,

where f is the fraction representing the natural abundance of the isotope and M is the mass number (weight) of the isotope.


The average atomic mass of the element is 49.237 u.

f1 = 0.3594 (35.94%)

M1 = 41.834 u

f2 = 1 - 0.3594 = 0.6406 (64.06%)

M2 = unknown


Hence,

The average atomic mass of the element = f1M1 + f2M2

49.237 = 0.3594 × 41.834 + 0.6406 × M2

34.202 = 0.6406 × M2

M2 = 53.390 u


Answer: The mass of the second isotope is 53.390 u.